# Non-planar ends are continuously unforgettable

### v1

@@ 5.1 Moving to the full mapping class group
By [20, Proposition 5.4.6], there exists $x\in E$ and a sequence $\{E_{n}\}_{n\in\mathbb{N}}$ of pairwise-homeomorphic and disjoint clopen subsets of $E$ such that $E\smallsetminus\{x\}=\bigcup_{n\in\mathbb{N}}E_{n}$. From this decomposition, for each $n\in\mathbb{N}$, we can find a involution $\tau_{n}\colon\thinspace E\to E$ such that $\tau_{n}(E_{n})=E_{n+1}$ and $\tau_{n}$ restricts to the identity on $E_{k}$ for $k\neq n,n+1$. Observe that $\lim_{n\to\infty}\tau_{n}$ is the identity. Moreover, [20, Corollary 5.2.3] implies that $\tau_{n}$ is a normal generator for $\operatorname{Homeo}(E)$. ∎
@@
The argument in the proof of Theorem 5.5 works for a larger class of 2-manifolds; however, it is not clear how to naturally define the class. But, for example, the proof also implies that if $M$ is an orientable infinite-genus 2-manifold with no planar ends and with $\operatorname{\mathcal{E}}(M)$ homeomorphic to $\omega+1$, then Theorem 5.5 holds for $M$. In this case, $\operatorname{Homeo}(\operatorname{\mathcal{E}}(M))$ is topologically isomorphic to $\mathrm{Sym}(\mathbb{N})$, and so contains a dense subgroup generated by involutions arbitrary close to the identity (namely, the subgroup consisting of finitely supported permutations).

### v2

@@ 5.1 Moving to the full mapping class group
By [21, Proposition 5.4.6], there exists $x\in E$ and a sequence $\{E_{n}\}_{n\in\mathbb{N}}$ of pairwise-homeomorphic and disjoint clopen subsets of $E$ such that $E\smallsetminus\{x\}=\bigcup_{n\in\mathbb{N}}E_{n}$. From this decomposition, for each $n\in\mathbb{N}$, we can find a involution $\tau_{n}\colon\thinspace E\to E$ such that $\tau_{n}(E_{n})=E_{n+1}$ and $\tau_{n}$ restricts to the identity on $E_{k}$ for $k\neq n,n+1$. Observe that $\lim_{n\to\infty}\tau_{n}$ is the identity. Moreover, [21, Corollary 5.2.3] implies that $\tau_{n}$ is a normal generator for $\operatorname{Homeo}(E)$. ∎
@@
The argument in the proof of Theorem 5.5 works for a larger class of 2-manifolds; however, it is not clear how to naturally define the class. For example, we can extend to another class of self-similar 2-manifolds using [10]. In particular, replacing the use of [21, Proposition 5.4.6] and [21, Corollary 5.2.3] in Lemma 5.7 with [10, Proposition 2.22] and [10, Theorem 3.12], respectively, then the following theorem has the same proof as Theorem 5.5.
Theorem 5.8.
Let $M$ and $M^{\prime}$ be orientable 2-manifolds. Suppose $M$ is infinite genus with no planar ends and that the end space of $M$ is countable with Cantor–Bendixson degree a successor ordinal and with Cantor–Bendixson rank one (i.e., it is homeomorphic to an ordinal of the form $\omega^{\alpha+1}+1$ equipped with its order topology, where $\omega$ is the first countable ordinal and $\alpha$ is a countable ordinal). If $G^{\prime}$ is a large subgroup of $\operatorname{MCG}(M^{\prime})$, then every continuous epimorphism from $\operatorname{MCG}(M)$ to $G^{\prime}$ is induced by a homeomorphism. ∎
We do not know if the proof of Theorem 5.5 can be extended to cover all infinite-genus self-similar 2-manifolds with no planar ends; however, we know the strategy cannot work for all infinite-genus 2-manifolds. For instance, consider the infinite-genus 2-manifold $M$ with no planar ends and whose end space $E$ is homeomorphic to $\omega\cdot 2+1$ (or equivalently, the two-point compactification $\widehat{\mathbb{Z}}$$\mathbb{Z}\cup\{\pm\infty\}$ of $\mathbb{Z}$). Then the closure of the finitely supported homeomorphisms in $\operatorname{Homeo}(E)$ of $E$, call it $G$, is an open neighborhood of the identity and a normal subgroup of $\operatorname{Homeo}(E)$ such that $\operatorname{Homeo}(E)/G$ is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{2}$ (this follows from the results in [10, Section 4]). In particular, $G$ is a neighborhood of the identity that cannot contain a normal generating set of $\operatorname{Homeo}(E)$, implying the techniques in the proof of Theorem 5.5 cannot be applied to $M$ (though, we note that this discussion provides enough information to make an argument that allows one to extend the conclusion of Theorem 5.5 to $\operatorname{MCG}(M)$).